# A survey of the monotonicity and non-contradiction of consensus methods and supertree methods

### v2

@@ 1 Introduction
In phylogenetics, consensus methods play a fundamental role concerning tree reconstruction and are therefore often used in biological studies (cf. [22, 7, 13]). For instance, when different genes of the same set of species lead to different gene trees or when different tree reconstruction methods come to different results, it may be hard to decide which of the given trees is the “true” tree in the sense of coinciding with the underlying (unknown) species tree. This is where consensus methods come into play — they use certain rules to summarize a set of trees to form a consensus tree. There are various such methods used in biology, which is why it makes sense to classify them according to their properties, some of which are desirable, while others might be undesirable. Several approaches in this regard have been made in the literature. Most prominently, Day and McMorris gave an overview of certain axioms and paradigms used to categorize consensus methods [8]. This axiomatic approach seems to have been inspired by the famous Arrovian result on weak orders [2], then considered in the area of partitions by Mirkin [21], before finally being considered in more tree-like contexts, outlined by McMorris and Powers in [19]. One popular axiom (e.g. discussed in [8, 19, 18]) is monotonicity, which we will re-visit in the present manuscript, along with a criterion which we will coin non-contradiction.
@@ 2.1 Basic phylogenetic concepts
We now first turn our attention to unrooted trees. A bipartition $\sigma$ of $X$ into two non-empty and disjoint subsets $A$ and $B$ is called an $X$-split (or split for short), and is denoted by $\sigma=A|B$. There is a natural relationship between $X$-splits and the edges of an unrooted phylogenetic $X$-tree $T$, because the removal of an edge $e$ induces such a bipartition of $X$. An $X$-split with $\min\{|A|,|B|\}=1$ is called trivial. In the following, the set of all induced (non-trivial) $X$-splits of $T$ will be denoted by $\Sigma(T)$ (or $\Sigma^{*}(T)$, respectively). Note that two $X$-splits $\sigma_{1}=A|B$ and $\sigma_{2}=\widetilde{A}|\widetilde{B}$ are called compatible if and only if at least one of the intersections $A\cap\widetilde{A}$, $A\cap\widetilde{B}$, $B\cap\widetilde{A}$ or $B\cap\widetilde{B}$ is empty. It is well-known (the so-called Buneman theorem or splits-equivalence theorem) that each set of compatible splits uniquely defines a tree, which can be found by the Tree Popping algorithm [20, 23]).
@@
and this often gets denoted by the shorthand $f(1)f(2)\ldots f(n)$. Moreover, we may assume without loss of generality that $f(1)=1$, because we consider characters equivalent if they correspond to the same $X$-split, even if the roles of 0 and 1 are interchanged. We call the set of binary characters induced by a phylogenetic $X$-tree with $B^{*}(T)$.
@@
In the rooted setting, instead of considering splits we consider clusters. A clade of a rooted phylogenetic tree $T$ is a pendant subtree $T^{\prime}$ of $T$, and a cluster is the set of leaves $Y\subseteq X$ of such a pendant subtree $T^{\prime}$ of $T$. Note that every vertex of a rooted phylogenetic tree $T$ induces a cluster as it can be regarded as the root of such a subtree of $T$. Clusters of sizes $1$ and $n$ are called trivial. We denote the set of all clusters of a rooted phylogenetic tree $T$ by $\mathcal{C}(T)$, and the set of its non-trivial clusters by $\mathcal{C}^{*}(T)$. Note that every cluster $Y\subset X$ induces an $X$-split, namely $\sigma=Y|(X\setminus Y)$. In the following, we call two clusters $c_{1},c_{2}$ compatible if the induced splits $\sigma_{1}$ and $\sigma_{2}$ of $c_{1}$ and $c_{2}$, respectively, are compatible and if additionally we have $c_{1}\subseteq c_{2}$ or $c_{2}\subseteq c_{1}$ or $c_{1}\cap c_{2}=\emptyset$.
@@ 2.2 Consensus methods
A consensus method is termed non-contradictory if each cluster (or split, respectively) of the output tree is compatible with at least one tree in the input profile.
@@
Next, we want to introduce two more classic consensus methods, namely the Adams and the Aho consensus. Our definitions are based on [6], but for further details, we refer the reader also to [5].
@@
In both cases, the hierarchy induced by the recursive partitioning can be used to reconstruct a unique rooted tree (which we refer to as $\varphi_{Ad}(\mathcal{P})$ or $\varphi_{Aho}(\mathcal{P})$, respectively) using the so-called BUILD algorithm [1]. Note, however, that both the Adams and the Aho consensus are only defined in the rooted setting and do not work for unrooted trees.
@@
By [6, p. 613], “All standard phylogenetic consensus methods (e.g., strict consensus, majority rule, loose consensus, and Adams consensus)”, are regular. It is easy to see that this, too, applies to Aho consensus.
@@ 2.3 MRP and MRC
The reason why it is not straightforward to regard supertree methods as consensus methods is that a supertree need not be unique. This problem can be overcome, e.g., by using an existing consensus method in order to summarize all supertrees into a single consensus tree [5]. Another argument against regarding supertree methods as consensus methods is that a supertree may contain splits (or clusters) that are not present in any of the input trees (even though this can also happen with some established consensus methods, see Figure 2: There, cluster $\{4,5\}$ is contained in the Adams and Aho consensus trees but in none of the input trees). However, in the present manuscript, we will not join this debate, but we will consider two established supertree methods as consensus methods in the case in which the respective supertree is unique and test them for refinement stability: Matrix Representation with Parsimony, or MRP for short, and Matrix Representation with Compatibility, or MRC for short. [^1] Note that – as we only consider cases where the output of MRP and MRC is unique – applying a regular consensus method like majority-rule to the output of MRP or MRC as suggested for instance in [5] would not change anything due to unanimity.
@@
Given a profile $\mathcal{P}=(T_{1},\ldots,T_{m})$ of phylogenetic $X$-trees, we consider the union of binary characters $\mathcal{B}(\mathcal{P}):=\bigcup\limits_{i=1}^{m}B^{*}(T_{i})$. Note that biologists often regard the multiset $\mathcal{B}(\mathcal{P})$ as a so-called alignment, i.e., as an ordered sequence of characters. However, as the order of characters does not play a role for the methods considered in the present manuscript, it is sufficient for us to consider the union. However, note that here we understand the union in the sense of a multiset, i.e., as two phylogenetic trees $T_{i}$ and $T_{j}$ might have a split in common, they might contribute the same character to $\mathcal{B}$, in which case both copies are kept. Now, an MP tree of $\mathcal{B}$ is called Matrix Representation with Parsimony tree of $\mathcal{P}$, or MRP tree of $\mathcal{P}$ for short. Again, the MRP tree need not be unique.
The second supertree method which we want to consider is Matrix Representation with Compatibility, or MRC for short. This method analyzes all characters of $B$ and finds a maximal compatible subset. Here, a set $S$ of binary characters is called compatible if all characters in $S$ are pairwise compatible, and two binary characters are called compatible if their corresponding splits are compatible. By the Buneman theorem, there is a unique tree $T$ corresponding to each such set of compatible characters, which is why we say that such a set induces $T$. Now, MRC works as follows: After a maximal compatible subset of characters in $B(\mathcal{P})$ is found and the corresponding tree is constructed, this tree is called Matrix Representation with Compatibility tree or MRC tree for short. Note that there may be more than one MRC tree.
Note that both MRP and MRC can only reconstruct unrooted trees as they deliver no information about the root position. So while we generally consider both the rooted and the unrooted setting in this manuscript, this does not hold true whenever we consider MRP or MRC.
@@ 3.1.1 Monotonicity
We prove the assertion concerning $\mathtt{loose}$ by providing an explicit counterexample. Let $X=\{1,2,3,4\}$ and $T_{1}=((1,2),3,4)$ be either rooted or unrooted, and let $S=(1,2,3,4)$ be the (rooted or unrooted) star-tree, i.e., the unique tree on $X$ with only one inner vertex. Let $\mathcal{P}=\{T_{1},S\}$. Then, for the loose consensus tree we have $\mathtt{loose}(\mathcal{P})=T_{1}$. However, if we refine $S$ to become $T_{2}=((2,3),1,4)$ and consider the profile $\mathcal{P}^{\prime}=(T_{1},T_{2})$, we get $\mathtt{loose}(\mathcal{P}^{\prime})=S$ (because the splits $12|34$ and $23|14$ are incompatible). So if we refine input tree $S$, we get a coarser consensus tree, namely $S$ instead of $T_{1}$, not a refined one. This completes the proof for $\mathtt{loose}$.
Concerning Adams and Aho consensus, we prove the statement by presenting an explicit counterexample. Let $T_{1}=(((1,2,3),4),5)$, $T_{2}=((1,2),(3,4),5)$ and $T_{2}^{\prime}=(((1,2),(3,4)),5)$, and let $\mathcal{P}=(T_{1},T_{2})$ and $\mathcal{P}^{\prime}=(T_{1},T_{2}^{\prime})$. Note that $T_{2}^{\prime}$ is a refinement of $T_{2}$, and hence any consensus method $\phi$ that is monotonic must have $\phi(P)\preceq\phi(P^{\prime})$. However, the Adams consensus tree of $\mathcal{P}$ is $\varphi_{Ad}(\mathcal{P})=T_{2}$, and the Adams consensus of $\mathcal{P}^{\prime}$ is tree $\varphi_{Ad}(\mathcal{P})^{\prime}=(((1,2),3,4),5)$. It follows that the Adams consensus method is not monotonic as $\varphi_{Ad}(\mathcal{P})\not\preceq\varphi_{Ad}(\mathcal{P}^{\prime})$.
@@
So in summary, of the established consensus methods, strict and majority-rule are monotonic, whereas the loose, Adams and Aho consensus methods are not. In Section 3.3 we will show that this also applies to MRP and MRC, unfortunately. But before we do so, we turn our attention to another property that consensus methods might have, namely that of being non-contradictory.

### v4

@@ 1 Introduction
In phylogenetics, consensus methods play a fundamental role concerning tree reconstruction and are therefore often used in biological studies (cf. [1, 2, 3]). For instance, when different genes of the same set of species lead to different gene trees or when different tree reconstruction methods come to different results, it may be hard to decide which of the given trees is the “true” tree in the sense of coinciding with the underlying (unknown) species tree. This is where consensus methods come into play — they use certain rules to summarize a set of trees to form a consensus tree. There are various such methods used in biology, which is why it makes sense to classify them according to their properties, some of which are desirable, while others might be undesirable. Several approaches in this regard have been made in the literature. Most prominently, Day and McMorris gave an overview of certain axioms and paradigms used to categorize consensus methods [4]. This axiomatic approach seems to have been inspired by the famous Arrovian result on weak orders [5], then considered in the area of partitions by Mirkin [6], before finally being considered in more tree-like contexts, outlined by McMorris and Powers in [7]. One popular axiom (e.g. discussed in [4, 7, 8]) is monotonicity, which was originally introduced in [8] as an example for a desirable property of consensus methods [4, Chapter 1.2] and which we will re-visit in the present manuscript, along with a new criterion which we will coin non-contradiction.
@@ 2.1 Basic phylogenetic concepts
We now first turn our attention to unrooted trees. A bipartition $\sigma$ of $X$ into two non-empty and disjoint subsets $A$ and $B$ is called an $X$-split (or split for short), and is denoted by $\sigma=A|B$. There is a natural relationship between $X$-splits and the edges of an unrooted phylogenetic $X$-tree $T$, because the removal of an edge $e$ induces such a bipartition of $X$. An $X$-split with $\min\{|A|,|B|\}=1$ is called trivial. In the following, the set of all induced (non-trivial) $X$-splits of $T$ will be denoted by $\Sigma(T)$ (or $\Sigma^{*}(T)$, respectively). Note that two $X$-splits $\sigma_{1}=A|B$ and $\sigma_{2}=\widetilde{A}|\widetilde{B}$ are called compatible if and only if at least one of the intersections $A\cap\widetilde{A}$, $A\cap\widetilde{B}$, $B\cap\widetilde{A}$ or $B\cap\widetilde{B}$ is empty. It is well-known (the so-called Buneman theorem or splits-equivalence theorem) that each set of pairwise compatible splits uniquely defines a tree, which can be found by the Tree Popping algorithm [14, 15].
@@
and this often gets denoted by the shorthand $f(1)f(2)\ldots f(n)$. Moreover, we may assume without loss of generality that $f(1)=1$, because we consider characters equivalent if they correspond to the same $X$-split, even if the roles of 0 and 1 are interchanged. We denote the set of non-trivial binary characters induced by a phylogenetic $X$-tree with $B^{*}(T)$.
@@
In the rooted setting, instead of considering splits we consider clusters. A clade of a rooted phylogenetic tree $T$ is a pendant subtree $T^{\prime}$ of $T$, and a cluster is the set of leaves $Y\subseteq X$ of such a pendant subtree $T^{\prime}$ of $T$. Note that every vertex of a rooted phylogenetic tree $T$ induces a cluster as it can be regarded as the root of such a subtree of $T$. Clusters of sizes $1$ and $n$ are called trivial. We denote the set of all clusters of a rooted phylogenetic tree $T$ by $\mathcal{C}(T)$, and the set of its non-trivial clusters by $\mathcal{C}^{*}(T)$. Note that every cluster $Y\subsetneq X$ induces an $X$-split, namely $\sigma=Y|(X\setminus Y)$ (note that if $Y=X$, $Y$ does not induce an $X$-split, which is why the condition $Y\subsetneq X$ is crucial). In the following, we call two clusters $c_{1},c_{2}$ compatible if we have $c_{1}\subseteq c_{2}$ or $c_{2}\subseteq c_{1}$ or $c_{1}\cap c_{2}=\emptyset$. Note that this necessarily implies that the splits induced by $c_{1}$ and $c_{2}$ are also compatible if and only if $c_{1}$ and $c_{2}$ are compatible.
@@ 2.2 Consensus methods
A consensus method is termed non-contradictory if each cluster (or split, respectively) of the output tree is compatible with at least one tree in the input profile, where a cluster (or split) is called compatible with a tree if it is pairwise compatible with all its induced clusters (or splits, respectively).
@@
Next, we want to introduce two more classic consensus methods, namely the Adams and the Aho consensus, which are both only defined for rooted trees. Our definitions are based on [9], but for further details, we refer the reader also to [20].
@@
In both cases, the hierarchy induced by the recursive partitioning can be used to reconstruct a unique rooted tree (which we refer to as $\varphi_{Ad}(\mathcal{P})$ or $\varphi_{Aho}(\mathcal{P})$, respectively) using the so-called BUILD algorithm [16].
@@
By [9, p. 613], “All standard phylogenetic consensus methods (e.g., strict consensus, majority rule, loose consensus, and Adams consensus)”, are regular. It is easy to see that this, too, applies to Aho consensus.
@@ 2.3 MRP and MRC
The reason why it is not straightforward to regard supertree methods as consensus methods is that a supertree need not be unique. This problem can be overcome, e.g., by using an existing consensus method in order to summarize all supertrees into a single consensus tree [20]. Another argument against regarding supertree methods as consensus methods is that a supertree may contain splits (or clusters) that are not present in any of the input trees (even though this can also happen with some established consensus methods, see Figure 2: There, cluster $\{4,5\}$ is contained in the Adams and Aho consensus trees but in none of the input trees). However, in the present manuscript, we will not join this debate, but we will consider two established supertree methods as consensus methods in the case in which the respective supertree is unique and test them for refinement stability: Matrix Representation with Parsimony, or MRP for short, and Matrix Representation with Compatibility, or MRC for short. [^1] Note that – as we only consider cases where the output of MRP and MRC is unique – applying a regular consensus method like majority-rule to the output of MRP or MRC as suggested for instance in [20] would not change anything due to unanimity. Note that as opposed to the consensus methods introduced above, MRP and MRC both are only defined for unrooted trees as they deliver no information about the root position.
@@
Given a profile $\mathcal{P}=(T_{1},\ldots,T_{m})$ of phylogenetic $X$-trees, we consider the union of binary characters $\mathcal{B}(\mathcal{P}):=\bigcup\limits_{i=1}^{m}B^{*}(T_{i})$. Note that biologists often regard the multiset $\mathcal{B}(\mathcal{P})$ as a so-called alignment, i.e., as an ordered sequence of characters. However, as the order of characters does not play a role for the methods considered in the present manuscript, it is sufficient for us to consider the union. However, note that here we understand the union in the sense of a multiset, i.e., as two phylogenetic trees $T_{i}$ and $T_{j}$ might have a split in common, they might contribute the same character to $\mathcal{B}$, in which case both copies are kept. Now, an MP tree of $\mathcal{B}$ is called Matrix Representation with Parsimony tree of $\mathcal{P}$, or MRP tree of $\mathcal{P}$ for short. Again, the MRP tree need not be unique. In particular, it can easily be seen that if $T$ is an MRP tree for some profile $\mathcal{P}$, all of its refinements are MRP trees for $\mathcal{P}$, too (because adding additional edges to a tree cannot increase its parsimony score).
The second supertree method which we want to consider is Matrix Representation with Compatibility, or MRC for short. This method analyzes all characters of $\mathcal{B}(\mathcal{P})$ and finds a maximum compatible subset. Here, a set $S$ of binary characters is called compatible if all characters in $S$ are pairwise compatible, and two binary characters are called compatible if their corresponding splits are compatible. By the Buneman theorem, there is a unique tree $T$ corresponding to each such set of compatible characters, which is why we say that such a set induces $T$. Now, MRC works as follows: After a maximum compatible subset of characters in $B(\mathcal{P})$ is found and the corresponding tree is constructed, this tree is called Matrix Representation with Compatibility tree or MRC tree for short. Note that there may be more than one MRC tree as more than one maximum set of compatible characters may exist in $\mathcal{B}(\mathcal{P})$.
@@ 3.1.1 Monotonicity
We prove the assertion concerning $\mathtt{loose}$ by providing an explicit counterexample. Let $X=\{1,2,3,4\}$ and $T_{1}=((1,2),3,4)$ be either rooted or unrooted, and let $S=(1,2,3,4)$ be the (rooted or unrooted) star-tree, i.e., the unique tree on $X$ with only one inner vertex. Let $\mathcal{P}=(T_{1},S)$. Then, for the loose consensus tree we have $\mathtt{loose}(\mathcal{P})=T_{1}$. However, if we refine $S$ to become $T_{2}=((2,3),1,4)$ and consider the profile $\mathcal{P}^{\prime}=(T_{1},T_{2})$, we get $\mathtt{loose}(\mathcal{P}^{\prime})=S$ (because the splits $12|34$ and $23|14$ are incompatible). So if we refine input tree $S$, we get a coarser consensus tree, namely $S$ instead of $T_{1}$, not a refined one. This completes the proof for $\mathtt{loose}$.
Concerning Adams and Aho consensus, we prove the statement by presenting an explicit counterexample. Let $T_{1}=(((1,2,3),4),5)$, $T_{2}=((1,2),(3,4),5)$ and $T_{2}^{\prime}=(((1,2),(3,4)),5)$, and let $\mathcal{P}=(T_{1},T_{2})$ and $\mathcal{P}^{\prime}=(T_{1},T_{2}^{\prime})$. Note that $T_{2}^{\prime}$ is a refinement of $T_{2}$, and hence any consensus method $\phi$ that is monotonic must have $\phi(\mathcal{P})\preceq\phi(\mathcal{P}^{\prime})$. However, the Adams consensus tree of $\mathcal{P}$ is $\varphi_{Ad}(\mathcal{P})=T_{2}$, and the Adams consensus of $\mathcal{P}^{\prime}$ is tree $\varphi_{Ad}(\mathcal{P})^{\prime}=(((1,2),3,4),5)$. It follows that the Adams consensus method is not monotonic as $\varphi_{Ad}(\mathcal{P})\not\preceq\varphi_{Ad}(\mathcal{P}^{\prime})$.
@@
While Proposition 3.2 can be seen as a negative result, there are cases when monotonicity can be guaranteed even for the loose, Adams and Aho consensus methods. This depends on certain properties of the input sets. We now analyze this further and start with the loose consensus method.
Proposition 3.3.
Let $\mathcal{P}=\{T_{1},\ldots,T_{k}\}$ and let $\mathcal{P}^{\prime}=\{T_{1}^{\prime},\ldots,T_{k}^{\prime}\}$ be such that $T_{i}\preceq T_{i}^{\prime}$ for all $i\in\{1,\ldots,k\}$. Moreover, let $\mathcal{P}^{\prime}$ be such that it permits a linear ordering, i.e., for all $i,j\in\{1,\ldots,k\}$ we either have $T_{i}^{\prime}\preceq T_{j}^{\prime}$ or $T_{j}^{\prime}\preceq T_{i}^{\prime}$. Let $\phi\in\{\mathtt{loose},\varphi_{Adams},\varphi_{Aho}\}$. Then we have: $\phi(\mathcal{P})\preceq\phi(\mathcal{P}^{\prime})$.
Proof.
Let $\mathcal{P}$ and $\mathcal{P}^{\prime}$ be as described in the proposition. Then, as $\mathcal{P}^{\prime}$ allows for a linear ordering, we may assume without loss of generality that $T_{i}^{\prime}\preceq T_{i+1}^{\prime}$ for all $i=1,\ldots,k-1$. Note that this implies $T_{k}^{\prime}$ is the most refined tree in $\mathcal{P}^{\prime}$, and as all splits of $T_{k}^{\prime}$ are necessarily compatible with all other trees in $\mathcal{P}^{\prime}$, we must have $\mathtt{loose}(\mathcal{P}^{\prime})=T_{k}^{\prime}$. Moreover, the Adams consensus as well as the Aho consensus for two trees $\widetilde{T}$ and $\widehat{T}$ with $\widetilde{T}\preceq\widehat{T}$ will always return $\widetilde{T}$, we have $\varphi_{Adams}(\mathcal{P}^{\prime})=\varphi_{Aho}(\mathcal{P}^{\prime})=T_{1}^{\prime}$, i.e., the coarsest tree in $\mathcal{P}^{\prime}$ is returned by these methods.
On the other hand, as we have $T_{i}\preceq T_{i}^{\prime}$ for all $i=1,\ldots,k$ by assumption, we know that the set $\Sigma(\mathcal{P})$ containing all splits occurring in any of the trees of $\mathcal{P}$ is a subset of $\Sigma(\mathcal{P}^{\prime})$, the set of splits of $\mathcal{P}^{\prime}$, i.e., $\Sigma(\mathcal{P})\subseteq\Sigma(\mathcal{P}^{\prime})$. By $T_{i}^{\prime}\preceq T_{k}^{\prime}$ for all $i=1,\ldots,k$, we know that $\Sigma(\mathcal{P}^{\prime})=\Sigma(T^{\prime}_{k})=\Sigma(\mathtt{loose}(\mathcal{P}^{\prime}))$. As by definition we must have $\Sigma(\mathtt{loose}(\mathcal{P}))\subseteq\Sigma(\mathcal{P})$, we conclude $\Sigma(\mathtt{loose}(\mathcal{P}))\subseteq\Sigma(\mathcal{P})\subseteq\Sigma(\mathcal{P}^{\prime})=\Sigma(\mathtt{loose}(\mathcal{P}^{\prime}))$, which shows that $\mathtt{loose}(\mathcal{P})\preceq\mathtt{loose}(\mathcal{P}^{\prime})$ as desired. Moreover, again as above, $\varphi_{Adams}(\mathcal{P})=\varphi_{Aho}(\mathcal{P})=T^{\star}$, where $T^{\star}$ is the coarsest tree in $\mathcal{P}$. Let $i^{\star}\in\{1,\ldots,k\}$ be such that $T_{i^{\star}}=T^{\star}$. Then, we have: $\varphi_{Adams}(\mathcal{P})=\varphi_{Aho}(\mathcal{P})=T^{\star}=T_{i^{\star}}\preceq T_{1}\preceq T_{1}^{\prime}=\varphi_{Adams}(\mathcal{P}^{\prime})=\varphi_{Aho}(\mathcal{P}^{\prime})$, which completes the proof. ∎
So in summary, of the established consensus methods, strict and majority-rule are monotonic, whereas the loose, Adams and Aho consensus methods are not, but the latter at least have the monotonicity property if the input profiles are such that the refined one allows for a linear ordering. In Section 3.2 we will show that these monotonicity statements also apply to MRP and MRC. But before we do so, we turn our attention to another property that consensus methods might have, namely that of being non-contradictory.